# Fingering instability in dewetting capillary nanosuspensions

### v1

@@ Abstract
This study investigates the fingering instability that forms when a capillary nanosuspension liquid bridge is stretched. The dewetting process is observed using a transparent lifted Hele-Shaw cell. The liquid bridge is stretched under constant acceleration, and the resulting instability patterns are recorded using two high-speed cameras. Finger-like structures, characteristic of the Saffman-Taylor instability are observed. The total length of the dendrites and the overlapped number of branches are quantified. We reveal the roles of microparticles, nanoparticles, and the secondary liquid during the fingering instability. Addition of microparticles to pure liquid enhanced finger length due to increased particle interactions and nucleation sites for bubbles. Addition of secondary fluid reduces fingering length by forming a strong interparticle network. Incorporation of Nanoparticles induces an early onset of cavitation and enhanced fingering instability. However, nanoparticles make the capillary suspensions’ overall micro-structure more homogeneous, reduce the sample variation in fingering patterns, and promote the even distribution of gel on both slides during splitting. These findings highlight the complex interactions governing dewetting in capillary (nano)suspensions. This knowledge has potential applications in microfluidics, 3D printing, and thin-film coatings, where controlling dewetting is crucial.
@@ 1 Introduction
While the fundamental behavior of liquid bridges has been well-characterized for simple systems, real-world applications often involve additional complexities, particularly the presence of particles. Understanding how particles interact with liquid bridges is crucial for applications ranging from materials processing to printing technologies. When microparticles are connected by the aqueous bridges, the oil is the bulk phase, substituting air in the following literature studies. These bridges play a crucial role in particle-surface interactions. When a particle impacts a wet surface, it first penetrates the liquid layer, generating elastic waves in both the particle and wall. These waves, along with potential plastic deformation depending on impact velocity, determine the initial particle rebound. During particle rebound, a liquid bridge forms, and the stresses in this stretching bridge determine the final particle velocity and potential substrate deposition [19]. The ratio between rebound and impact velocities, defined as the restitution coefficient, characterizes the energy dissipation during these collisions [20]. Grohn et al. [21] measured the restitution coefficient and bridge breakage for varying liquid volumes, finding that more liquid leads to longer rupture duration. Additional studies [22, 23] showed that in oblique impacts, wet particles exhibit higher tangential restitution coefficients than dry ones due to lubrication effects. The model presented by Davis [24] extended this to multi-wet body collisions, predicting outcomes from bouncing to clumping based on the interplay of inertia, viscosity, and surface tension forces.
@@
The complexity of liquid bridge behavior increases substantially when both micro and nanoparticles are present in the system. Based on our previous studies, we discovered that capillary nanosuspensions are normal capillary suspensions with a low volume fraction of nanoparticles inside the secondary liquid. In ceramic systems, these nanoparticles can strengthen sintering necks and form dense bridges [1, 4]. However, in wet precursor systems, nanoparticle addition decreases the yield stress, enhances interparticle movement, and promotes secondary liquid distribution, making the network more compressible and homogeneous [1, 4, 25].
@@
This study investigates the dynamic behavior of capillary nanosuspensions under the combined influence of thin-gap stretching and high acceleration. We employ a novel experimental setup, incorporating high-speed imaging, to directly observe the collective dewetting process. Our findings provide crucial insights into the dewetting dynamics of capillary nanosuspensions, with potential implications for controlling their rheological properties and optimizing their use in applications such as 3D printing and thin-film fabrication. The dynamic stretching of capillary nanosuspensions can lead to complex dewetting phenomena, including the formation of fingering instabilities. These instabilities arise from the interplay of capillary forces, viscous forces, and the presence of particles, making their accurate prediction a significant challenge. During fast stretching, the mass balance is not easily determinable even for Newtonian fluids, as the plates remain wetted by a thin residual liquid layer when the meniscus recedes [7]. The complexity of the system and high acceleration bring challenges in modeling the fingering instability mathematically, meanwhile, this work focuses on the fingering number, total dendritic length changes, and cavitation effect caused by additional inter-particle bridges and nanoparticles. The findings show the complex interplay between the properties of the microparticles, nanoparticles, liquid bridges, and instabilities in the flow within the confined gap.
@@ 2.2 Experimental setup
Experiments performed within this study are based on a rapidly stretched liquid bridge forming between two horizontally aligned substrates. The stretching process is induced by the motion of the lower substrate with the upper substrate remaining fixed. A sketch of the experimental setup is shown in Figure 1.
Figure 1: Schematic configuration of the liquid bridge stretching experiment (1) High-speed camera, (2) High Power LED, (3) Navitar Zoom Objective, (4) Gel/Liquid placed between an upper and lower substrate, (5) Linear actuator, 6) Telecentric objective, (7) High-speed camera (8) Confocal sensor. The scale bars on the exemplary images are 5 mm.
@@
Two separate high-speed video systems are utilized for observation. The first high-speed camera (Phantom T1340, labeled “1”) with a pixel size of 27 $\times$ 27 $\mathrm{\mu}$m records the liquid bridge from the top through the transparent upper substrate. A frame rate of 1000 fps and a resolution of 1024 px $\times$ 976 px is chosen for the experiments with a spatial resolution of 50 $\mathrm{\mu}$m/px. Illumination and magnification are achieved using a custom-built microscope comprising an optical tube (InfiniTube Special, Infinity Photo-Optical) illuminated by a 7 W high-power green LED ($\lambda\approx 532$ nm, ILA iLA.LPS v3, labeled as “7” in Fig. 1a). A dichroic mirror and two bandpass filters, mounted on a filter cube (Thorlabs DFM1/M, labeled as “9”), direct the excitation light through the objective and filter the fluorescence signal emitted by the excited particles. The system employs an Infinity Photo-Optical IF-3 objective (“3”) with 1$\times$ magnification. The images generated with a high-speed camera (“1”) are referred to as the top view in the following. The second high-speed camera (Phantom Miro Lab 110, labeled “7”) with a pixel size of 20 $\times$ 20 $\mathrm{\mu}$m records the side profile using a telecentric objective (“6”) and a high Power LED (“2”), the images generated with this camera are hereafter referred to as side view. The camera is operated with a frame rate of 3200 fps and a resolution of 800 $\times$ 640 px with a spatial resolution of 50 $\mathrm{\mu}$m/px. The distance between the substrates is measured with a confocal displacement sensor (Micro Epsilon IFS2407-0.1, “8”) with an accuracy of up to 3 nm within a maximum measuring range of 0.1 mm. The gap of the experiments in this study is set to 50 $\mathrm{\mu}$m.
@@
The constancy of sample loading is of great importance in our experiments. For normal suspensions with no liquid bridges and pure particles, we manually stirred the sample with a spatula before 50 $\mathrm{\mu}$l volume was taken by the syringe and dosed onto the bottom plate. For capillary (nano)suspensions, the exact dosing is impossible, since the compression forces can expel the bulk liquid and increase the volume fraction. The gels were manually scooped out and spread onto the bottom plate without being compressed extensively.
@@ 2.3 Image processing
Figure 2: (a-c) Exemplary images of the top view dendritic classification method. (d) Dendritic skeleton length and overlapped branches with the region of interest (ROI) over $R$, the increasing radius of ROI.
After training Ilastik to differentiate fingers from the background, the processed images (Fig. 2b) were then binarized and skeletonized using a self-written MATLAB code. Starting from the center of mass of the first processed image (Fig. 2c), a region of interest (ROI) was defined and expanded with each increment in a ring radius ($R$) of 5 pixels (250 $\mathrm{\mu}$m). This 5-pixel increment was determined experimentally to reduce noise while preserving valuable information. The pixels within the skeletons were tallied as the ring radius increased, serving as an indicator of the total area of the branches, named as dendritic skeleton length (50 $\mathrm{\mu}$m/pixel). Analysis and plotting of the images extended from the first to the tenth frame, and branches intersecting the ring peripheries were also calculated. Due to the cavitation effect and fingering instability, the frame chosen for analysis was where the cavitation stabilized and voids emerged. If no cavitation was observed, the frame with the highest total dendritic length was chosen. This method simplifies branches of different sizes into skeletons with unity pixels, providing a representative measure of dendritic skeleton length.
@@ 3.1 Suspension dewetting
Figure 3: Three suspensions of (a) small initial diameter 12 mm, and larger initial diameter 36 mm where (b) fingering instability or (c) cavitation is dominant, at different time stamps, dewetted with an acceleration of 50 m/s2. The scale bars are 5 mm.
the fingering pattern for a low sample volume (12 $\mathrm{\mu}$m initial diameter) resembles that of a Newtonian fluid under high substrate acceleration, as observed by Brulin et al. [12]. Dewetting begins at $t=5$ ms, the liquid begins to dewet the top plate, and the lost volume can only be provided by the outer due to volume conservation. Intensified fluorescent signals indicate particle accumulation at the fingering edges. At $t=7$ ms, the dendritic branches elongate vertically, causing fluid to accumulate along the vertical direction of the branches. This accumulation increases the concentration of particles in the vertical region, leading to a more pronounced fluorescence signal.
During dewetting, for particles to be suspended and follow the flow, the lift force $F_{L}$ must exceed the gravitational force $F_{G}$. The gravitational force $F_{G}=\frac{1}{6}\pi d_{p}^{3}\Delta\rho$, where $d_{p}$ particle diameter and $\Delta\rho$ is the density difference. The lift force for particles near the plate can be expressed as seen in Eq. 1 [27],
@@
Affirming that the particles do follow the liquid flow, we can focus on the intensity of fingering patterns. At $t=7$ ms, branches are brighter at the edges. The change in the position of the discontinued bright spots indicates the movement of microparticles at $t=10$ ms. At $t=16$ ms, the particles migrate towards the center region, forming a particle-laden major bridge, as shown in Fig 4b.
Figure 4: The total dendritic skeleton length and overlapped branches with ROI of suspensions of (a) small initial diameter 12 mm, and larger initial diameter 36 mm where (b) fingering instability or (c) cavitation is dominant, dewetted with an acceleration of 50 m/s2, where instability or cavitation is dominant. The data are plotted over the normalized radius. Side view of the suspension samples where (b) instability and (c) cavitation dominant, at $t=14$ ms. The scale bars are 5 mm.
@@
For higher sample volumes with an initial diameter of 36 mm (Fig. 3b and c), the dewetting patterns differ significantly. The local stress differences caused by fast flow can either be compensated by instability and radius decreasing or cavitation bubble appearance and growth [28]. In both Fig. 3b and 3c, cavitation begins within the initial 5 ms of stretching, typically driven by the presence of foreign particles decreasing the total free energy needed for bubble nucleation [29, 30, 31]. It was observed that the cavitation formation appears almost simultaneously, within a sub-microsecond. In Fig. 3b, the Saffman–Taylor instability dominates with the minimum effect of cavitation, and the branches are mostly interconnected ($t=10$ ms). For the sample shown in Fig. 3c ($t=5$ ms), the cavitation begins with the appearance of instability. As cavitation and instability interface collide, the connected networks break.
@@
The experimental values for the total dendritic skeleton length and the overlapped branch number (skeletonized branch number values on ROI) between instability-dominated (Fig. 3b) and cavitation-dominated (Fig. 3c) samples are shown in Fig. 4a. The instability-dominated case shows a nearly 40 % increase in branch length and a 100 % increase in overlapped branch number. This phenomenon agrees with the estimation of Brulin et al. [7], where the cavitation lowers the branch number significantly (up to 120%). The broken branches caused by cavitation and instability eventually retract, leading to small particle-laden bridges that appear as single dots in Fig. 3c.
3.2 Capillary suspensions dewetting

### v2

# Dewetting Fingering Instability in Capillary Suspensions: Role of Particles and Liquid Bridges

@@ Abstract
This study investigates the fingering instability that forms during stretching of capillary suspensions with and without added nanoparticles. The dewetting process is observed using a transparent lifted Hele-Shaw cell. The liquid bridge is stretched under constant acceleration, and the resulting instability patterns are recorded using two high-speed cameras. Finger-like structures, characteristic of the Saffman-Taylor instability are observed. The total length of the dendrites and the intersecting number of branches are quantified. We reveal the roles of microparticles, nanoparticles, and the secondary liquid during the fingering instability. The addition of microparticles to pure liquid enhanced finger length due to increased particle interactions and nucleation sites for bubbles. The addition of secondary fluid reduces fingering length by forming a strong interparticle network. Incorporation of nanoparticles induces an early onset of cavitation and enhances fingering instability. However, nanoparticles make the capillary suspensions’ overall microstructure more homogeneous, reduce the sample variation in fingering patterns, and promote the even distribution of gel on both slides during splitting. These findings highlight the complex interactions governing dewetting in capillary (nano)suspensions. This knowledge has potential applications in microfluidics, 3D printing, and thin-film coatings, where controlling dewetting is crucial.
@@ 1 Introduction
In addition to interfacial instabilities, the rapid stretching of confined liquid bridges can trigger cavitation, which has distinct characteristics in a Hele-Shaw cell geometry compared to bulk systems. Due to the effects of confinement, bubbles preferentially expand in the planar direction rather than spherically. With decreasing gap height, the bubble lifetime increases by up to 50 % while the maximum projected bubble radius remains constant [19]. At high separation velocities, cavitation appears and the force response shows a plateau during which bubbles grow. When the bubbles eventually break, outside air rapidly enters the cavity, causing an abrupt force drop [20]. The presence of microscopic nuclei can lead to heterogeneous nucleation at pressures exceeding the theoretical cavitation threshold [21]. This interplay between cavitation and fingering becomes particularly complex in Hele-Shew cells, as both phenomena compete for stress relaxation during stretching via fingering through edge effects and cavitation through bulk deformation.
@@
While the fundamental behavior of liquid bridges has been well-characterized for simple systems, real-world applications often involve additional complexities such as additional particles. In capillary suspensions, microparticles form a sample-spanning network by aqueous bridge connections at the micrometer scale, with oil as the bulk phase rather than air. Understanding these particle-liquid bridge interactions is crucial for applications ranging from materials processing to printing technologies and the dynamics of particle-liquid bridge interactions have been extensively studied at the single-particle scale using millimeter-sized particles. When a particle impacts a wet surface, it first penetrates the liquid layer, generating elastic waves in both the particle and wall. These waves, along with potential plastic deformation at higher impact velocities, influence the initial particle rebound. During rebound, a liquid bridge forms, and the stresses in this stretching bridge determine the final particle velocity and potential substrate deposition [22]. The ratio between rebound and impact velocities, defined as the restitution coefficient, characterizes the energy dissipation during these collisions [23]. Grohn et al. [24] measured the restitution coefficient and bridge breakage for varying liquid volumes, finding that more liquid leads to longer rupture duration. Additional studies [25, 26] showed that in oblique impacts, wet particles exhibit higher tangential restitution coefficients than dry ones due to lubrication effects. The model presented by Davis [27] extended this to multi-wet body collisions, predicting outcomes from bouncing to clumping based on the interplay of inertia, viscosity, and surface tension forces.
@@
The complexity of liquid bridge behavior increases substantially when both micro and nanoparticles are present in the system. Based on our previous studies, we discovered that capillary nanosuspensions are normal capillary suspensions with a low volume fraction of nanoparticles inside the secondary liquid. In ceramic systems, these nanoparticles can strengthen sintering necks and form dense bridges [1, 4]. However, in wet precursor systems, nanoparticle addition decreases the yield stress, enhances interparticle movement, and promotes secondary liquid distribution, making the network more compressible and homogeneous [1, 4, 28, 29].
@@
This study investigates the dynamic behavior of capillary nanosuspensions under the combined influence of thin-gap stretching and high acceleration. We employ a novel experimental setup, incorporating high-speed imaging, to directly observe the collective dewetting process. Our findings provide crucial insights into the dewetting dynamics of capillary nanosuspensions, with potential implications for controlling their rheological properties and optimizing their use in applications such as 3D printing and thin-film fabrication. The dynamic stretching of capillary nanosuspensions can lead to complex dewetting phenomena, including the formation of fingering instabilities. These instabilities arise from the interplay of capillary forces, viscous forces, and the presence of particles, making their accurate prediction a significant challenge. During fast stretching, the mass balance is not easily determinable even for Newtonian fluids, as the plates remain wetted by a thin residual liquid layer when the meniscus recedes [7]. The complexity of the system and high acceleration bring challenges in modeling the fingering instability mathematically, meanwhile, this work focuses on the fingering number, total dendritic length changes, and cavitation effect caused by additional interparticle bridges and nanoparticles. The findings show the complex interplay between the properties of the microparticles, nanoparticles, liquid bridges, and instabilities in the flow within the confined gap.
@@ 2.2 Experimental setup
Experiments performed within this study are based on a rapidly stretched liquid bridge forming between two horizontally aligned substrates. The stretching process is induced by the motion of the lower substrate with the upper substrate remaining fixed. A sketch of the experimental setup is shown in Fig. 1.
Figure 1: Schematic configuration of the liquid bridge stretching experiment (1) High-speed camera, (2) High Power LED, (3) Navitar Zoom Objective, (4) Gel/Liquid placed between an upper and lower substrate, (5) Linear actuator, (6) Telecentric objective, (7) High-speed camera (8) Confocal sensor (9) Telecentric lighting. The scale bars on the exemplary images are 5 mm.
@@
Two separate high-speed video systems are utilized for observation. The first high-speed camera (Phantom T1340, labeled “1”) with a pixel size of 27 $\times$ 27 $\mathrm{\mu}$m records the liquid bridge from the top through the transparent upper substrate, where the bright pixels represent the sample. A frame rate of 1000 fps and a resolution of 1024 px $\times$ 976 px is chosen for the experiments with a spatial resolution of 50 $\mathrm{\mu}$m/px. Illumination and magnification are achieved using a custom-built microscope comprising an optical tube (InfiniTube Special, Infinity Photo-Optical) illuminated by a 7 W high-power green LED ($\lambda\approx 532$ nm, ILA iLA.LPS v3, labeled as “2”). A dichroic mirror and two bandpass filters, mounted on a filter cube (Thorlabs DFM1/M, labeled as “9”), direct the excitation light through the objective and filter the fluorescence signal emitted by the excited particles. The system employs an Infinity Photo-Optical IF-3 objective (“3”) with 1$\times$ magnification. The images generated with a high-speed camera (“1”) are referred to as the top view in the following. The second high-speed camera (Phantom Miro Lab 110, labeled “7”) with a pixel size of 20 $\times$ 20 $\mathrm{\mu}$m records the side profile using a telecentric objective (“6”) and a telecentric lighting (“9”), the images generated with this camera are hereafter referred to as side view, where the bright pixels are the background and the dark pixels are the substrates and samples in between. The camera is operated with a frame rate of 3200 fps and a resolution of 800 $\times$ 640 px with a spatial resolution of 50 $\mathrm{\mu}$m/px. The distance between the substrates is measured with a confocal displacement sensor (Micro Epsilon IFS2407-0.1, “8”) with an accuracy of up to 3 nm within a maximum measuring range of 0.1 mm. The gap of the experiments in this study is set to 50 $\mathrm{\mu}$m.
@@
The constancy of sample loading is of great importance in our experiments. For normal suspensions with no liquid bridges and pure particles, we manually stirred the sample with a spatula before 50 $\mathrm{\mu}$l volume was taken by the syringe and dosed onto the bottom plate. For capillary (nano)suspensions, the exact dosing is impossible, since the compression forces can expel the bulk liquid and increase the volume fraction. The gels were manually scooped out and spread onto the bottom plate, then the bottom plate was slowly brought to the desired position/gap without compressing the samples extensively.
@@ 2.3 Image processing
Figure 2: (a-c) Exemplary images of the top view dendritic classification method. (d) Dendritic skeleton length and intersecting branches with the region of interest (ROI) over $R$, the increasing radius of ROI, the intersection of branches with ROI are shown in yellow.
After training Ilastik to differentiate fingers from the background, the processed images (Fig. 2b) were then binarized and skeletonized using a self-written MATLAB code. Starting from the center of mass of the first processed image (Fig. 2c), a region of interest (ROI) was defined and expanded with each increment in a ring radius ($R$) of 5 pixels (250 $\mathrm{\mu}$m). This 5-pixel increment was determined experimentally to reduce noise while preserving valuable information. The pixels within the skeletons were tallied as the ring radius increased, serving as an indicator of the total area of the branches, named as dendritic skeleton length (50 $\mathrm{\mu}$m/pixel). Analysis and plotting of the images extended from the first to the tenth frame, the total dendritic skeleton lengths are plotted. Besides, the numbers of intersecting branches with ROI are also shown as the numbers of skeletons intersecting with the ROI ring peripheries, as shown by the yellow dots in Fig. 2c.
Due to the diverse pattern formation mechanisms observed, the frame selection criteria were adapted to the dominant behavior in each sample. For cavitation-dominant samples, where multiple disconnected void structures emerge, we analyzed the frame where cavitation had stabilized and distinct voids were visible. In contrast, for instability-dominant samples, which typically show connected dendritic patterns converging toward the center, we selected the frame exhibiting the maximum total dendritic length. This skeletonization method transforms branches of varying thicknesses into unity-pixel skeletons, providing a standardized measure of pattern length regardless of branch thickness. This approach accounts for both connected and isolated segments, as the concept of a main skeleton is primarily applicable to instability-dominated samples and not meaningful for cavitation-dominated cases where multiple independent structures form.
@@ 3.1 Suspension dewetting
Figure 3: Three suspensions of (a) small initial diameter 12 mm, and larger initial diameter 36 mm where (b) fingering instability or (c) cavitation is dominant, at different timestamps, dewetted with an acceleration of 50 m/s2. The gap height at each timestamp is $H$. The scale bars are 5 mm.
the fingering pattern for a low sample volume (12 mm initial diameter) resembles that of a Newtonian fluid under high substrate acceleration, as observed by Brulin et al. [12]. Dewetting begins at $t=5$ ms, the liquid begins to dewet the top plate, and the lost volume can only be provided by the outer due to volume conservation. Intensified fluorescent signals indicate particle accumulation at the fingering edges. At $t=7$ ms, the dendritic branches elongate vertically, causing fluid to accumulate along the vertical direction of the branches. This accumulation increases the concentration of particles in the vertical region, leading to a more pronounced fluorescence signal.
During dewetting, for particles to be suspended and follow the flow, the lift force $F_{L}$ must exceed the gravitational force $F_{G}$. The gravitational force $F_{G}=\frac{1}{6}\pi d_{p}^{3}\Delta\rho$, where $d_{p}$ is the particle diameter and $\Delta\rho$ is the density difference. The lift force for particles near the plate can be expressed as seen in Eq. 1 [31],
@@
Affirming that the flow velocity exceeding the critical velocity particles require to suspend in the liquid, the intensity of fingering patterns can be used to identify the microparticle displacement. At $t=7$ ms, branches are brighter at the edges. The change in the position of the discontinued bright spots indicates the movement of fluorescently labeled microparticles from edges towards the knots at $t=10$ ms. At $t=14$ 16 ms, the intensified signal toward the middle region indicates the directional particles migration, forming a particle-laden major filament.
@@
For higher sample volumes with an initial diameter of 36 mm (Fig. 3b and c), the dewetting patterns differ significantly. The local stress differences caused by fast flow can either be compensated by instability and radius decreasing or cavitation bubble appearance and growth [20]. In both Fig. 3b and 3c, cavitation begins within the initial 5 ms of stretching, typically driven by the presence of foreign particles decreasing the total free energy needed for bubble nucleation [21, 32, 33]. It was observed that the cavitation formation appears almost simultaneously, within a submicrosecond. In Fig. 3b, the Saffman–Taylor instability dominates with the minimum effect of cavitation, and the branches are mostly interconnected ($t=10$ ms). For the sample shown in Fig. 3c ($t=5$ ms), the cavitation begins with the appearance of instability. As cavitation and instability interface collide, the connected networks break. One interesting observed phenomenon is that the branches start to widen between $t=10$ 16 ms, this is especially visible for Fig. 3a and  3b. At the initial stage, the thickness of fingering instability branches is primarily driven by the extreme confinement effect of the small gap height ($\lambda=H_{0}/D_{0}<1$). Due to the small initial heights between plates, the liquids flow in a quasi-2D plane and the shear near the wall is maximum, this results in flows dominated by viscous forces and high pressure gradient caused thin fingers (with vertical filaments) connecting both plates [20]. After the filaments break in the middle due to the stretching, surface tension causes the remaining liquid to retract and redistribute on each plate, leading to an apparent widening of the branches as the fluid relaxes and spreads laterally.
The experimental values for the total dendritic skeleton length and the intersecting branch number (skeletonized branch number values on ROI) between instability-dominated (Fig. 3b) and cavitation-dominated (Fig. 3c) samples are shown in Fig. 4a.
Figure 4: Total dendritic skeleton length and intersecting branches with ROI of suspensions of larger initial diameters 36 mm where dewetted with an acceleration of 50 m/s2. Their side views are as shown, where in (b) fingering instability dominates and in (c) cavitation dominates. The scale bars are 5 mm.
@@
The instability-dominated case shows a nearly 40 % increase in branch length and a 100 % increase in intersecting branch number. This phenomenon agrees with the estimation of Brulin et al. [7], where the cavitation lowers the branch number significantly (up to 120%). The broken branches caused by cavitation and instability eventually retract, leading to small particle-laden bridges that appear as single dots in Fig. 3c. Their side view differences shown in Fig 4b and 4c, where instability dominant pattern retracts inward and forms limited amount of thick filaments. Meanwhile the cavitation-dominanted pattern spreads outward due to the extensive amount of freshly nucleated bubbles to form multiple single filaments during stretching. The transition between domination types was unable to be clearly predicted but likely to be associated with microparticle assembly due to compression, inducing the onset of cavitation.
3.2 Capillary suspension dewetting